# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@ 2.3. Fourier–Mukai transform
$\overline{\mathbb{FM}}:\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{-})\to\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{+})$
denote the Fourier–Mukai transform in this case, where $\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})$ denotes the completion of $K^{0}_{\mathbb{T}}(\overline{X}_{\pm})$ along the augmentation ideal (see [EG00]).
By definition of the $\mathbb{T}$ action on $B$, $E$, and $F$, the (completion of the) equivariant cohomology of $B$ is given by
$\hat{H}^{*}_{\mathbb{T}}(B)=H^{*}(B)[\![z_{i},x_{i}]\!]_{i=1}^{n}$
where recall that $x_{i}=c_{1}^{\mathbb{T}}(L_{i}^{\vee})$ and $z_{i}=c_{1}^{\mathbb{T}}(M_{i}^{\vee})$. The equivariant line bundles $L_{i}\to B$ and $M_{i}\to B$ for $1\leq i\leq n$ determine a map
$v:[B/(\mathbb{C}^{*})^{2n}]\to[\{pt\}/(\mathbb{C}^{*})^{2n}].$
@@
${\left[X_{\pm}/(\mathbb{C}^{*})^{2n}\right]}$${\left[\overline{X}_{\pm}/(\mathbb{C}^{*})^{2n}\right]}$${\left[B/(\mathbb{C}^{*})^{2n}\right]}$[${\left[\{pt\}/(\mathbb{C}^{*})^{2n}\right].}$$\scriptstyle{\pi_{\pm}}$$\scriptstyle{\tilde{v}_{\pm}}$$\scriptstyle{\overline{\pi}_{\pm}}$$\scriptstyle{v}$
@@
with $H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{\pm},$ we have the identification
${\pi_{\pm}}_{*}=\operatorname{id}_{B}\otimes_{\mathbb{C}}{\overline{\pi}_{\pm}}_{*}$
@@
$\hat{\operatorname{ch}}^{\mathbb{T}}:\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})\to\hat{H}^{*}_{\mathbb{T}}(\overline{X}_{\pm})$
is an isomorphism (see [EG00]; here we use $\hat{\operatorname{ch}}^{\mathbb{T}}$ to denote the induced map defined on the completion $\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{\pm})$). Because it is induced by a derived equivalence, the map $\overline{\mathbb{FM}}$ is also an isomorphism. Proposition 2.9 implies that $\overline{\mathbb{U}}_{H}:=\hat{\operatorname{ch}}^{\mathbb{T}}\circ\overline{\mathbb{FM}}\circ(\hat{\operatorname{ch}}^{\mathbb{T}})^{-1}$ is also an isomorphism, and may be defined before localizing with respect to $\mathbb{T}$. In other words, $\overline{\mathbb{U}}_{H}$ gives a well-defined map
$\overline{\mathbb{U}}_{H}:\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{-}\to\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{+}.$
Again using the identification $\hat{H}^{*}_{\mathbb{T}}(X_{\pm})=H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\ldots,\sigma_{r}]/\hat{I}_{\pm}$ together with the explicit formula for $\mathbb{U}_{H}$ in (2.3.5), we also see that $\mathbb{U}_{H}=\operatorname{id}_{B}\otimes_{\mathbb{C}}\overline{\mathbb{U}}_{H}$ which implies the first statement.
Given $\overline{\alpha}\in\hat{H}^{*}_{\mathbb{T}}(\overline{X}_{-})$, let $A=(\operatorname{ch}^{\mathbb{T}})^{-1}(\overline{\alpha})\in\hat{K}^{0}_{\mathbb{T}}(\overline{X}_{-})$. Applying the equivariant Hirzebruch–Riemann–Roch theorem (see (3.1.2) below) twice, we see that for $A\in\hat{K}^{0}_{\mathbb{T}}(X_{-})$,
(2.3.7) ${\overline{\pi}_{-}}_{*}\left(\overline{\alpha}\operatorname{Td}_{\overline{X}_{-}}\right)$ $=\chi(A)$
$=\chi(\overline{\mathbb{FM}}(A))$
$={\overline{\pi}_{+}}_{*}\left(\operatorname{ch}^{\mathbb{T}}(\overline{\mathbb{FM}}(A))\operatorname{Td}_{\overline{X}_{+}}\right)$
$={\overline{\pi}_{+}}_{*}\left(\mathbb{U}_{H}(\overline{\alpha})\operatorname{Td}_{\overline{X}_{+}}\right).$
The pullback $\tilde{v}_{\pm}^{*}(\operatorname{Td}_{\overline{X}_{\pm}})$ is the relative Todd class $\operatorname{Td}_{X_{\pm}}$. This fact together with (2.3.7) allows us to upgrade (2.3.6) to the following commutative diagram:
@@
3. Gromov–Witten theory and $I$-functions
Let $X$ be a smooth semi-projective variety. In what follows, $X$ will also be equipped with an action by a torus $\mathbb{T}$. Following [CIJS15], we assume that all $\mathbb{T}$-weights appearing in the representation $H^{0}(X,\mathcal{O}_{X})$ lie in a strictly convex cone, and that $H^{0}(X,\mathcal{O}_{X})^{\mathbb{T}}=\mathbb{C}.$ This implies in particular that $X^{\mathbb{T}}$ is compact.
@@ 3. Gromov–Witten theory and $I$-functions
$\int_{X}\gamma:=\sum_{F}\int_{F}\frac{(\gamma)|_{F}}{e_{\mathbb{T}}(N_{F|X})}\in S_{\mathbb{T}}$
@@
$\left\langle\alpha,\beta\right\rangle_{X}:=\sum_{F}\int_{F}\frac{(\alpha\cup\beta)|_{F}}{e_{\mathbb{T}}(N_{F|X})}.$
@@
$R[\![Q]\!]=\left\{\sum_{d\in\operatorname{Eff}(X)}a_{d}Q^{d}\mid a_{d}\in R\right\}$
@@
$\mathcal{H}_{X}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}(\!(z^{-1})\!)[\![Q]\!]$
@@
(3.0.1) $\Omega(f,g)=\operatorname{Res}_{z=0}\left\langle f(-z),g(z)\right\rangle_{X}dz.$
@@
with $\mathcal{H}_{X}^{+}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}[z][\![Q]\!]$ and $\mathcal{H}_{X}^{-}=z^{-1}H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}[\![z^{-1}]\!][\![Q]\!]$.
For $\alpha_{1},\ldots,\alpha_{m}\in H^{*}_{\mathbb{T}}(X)$, $k_{1},\ldots,k_{m}\geq 0$, and $d\in\operatorname{Eff}(X)$, define the degree-$d$ genus-0 Gromov–Witten invariant of $X$:
$\left\langle\alpha_{1}\psi_{1}^{k_{1}},\ldots,\alpha_{m}\psi_{m}^{k_{m}}\right\rangle_{0,m,d}^{X}=\int_{[\mathcal{M}_{0,m}(X,d)]^{vir}}\prod_{i=1}^{m}ev_{i}^{*}(\alpha_{i})\psi_{i}^{k_{i}}.$
If $X$ is not compact, the integral on the right hand side is defined after localization.
@@
(3.0.2) $\mathcal{F}_{X}({\bm{t}}(z))=\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle{\bm{t}}(\psi_{1}),\ldots,{\bm{t}}(\psi_{m})\right\rangle_{0,m,d}^{X},$
@@
$\mathcal{L}_{X}=\{(p,q)\in T^{*}\mathcal{H}_{X}^{+}\mid p=d_{q}\mathcal{F}_{X}\}.$
@@
$-z1+{\bm{t}}(z)+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},{\bm{t}}(\psi_{2}),\ldots,{\bm{t}}(\psi_{m+1})\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
@@
$J^{X}(t,z)=-z1+t+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\sum_{i\in I}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},t,\ldots,t\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
@@
$J^{X}(\tau,z)+\sum_{i\in I}C^{i}(z)\frac{\partial}{\partial{t^{i}}}J^{X}(t,z)|_{t=\tau}$
@@
3.1. Integral lattice
@@
$\mu(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X)}{2}\right)\phi,$
@@
Secondly, we define $\deg_{0}:H^{*}_{\mathbb{T}}(X)\to H^{*}_{\mathbb{T}}(X)$ be the degree operator, defined by
@@
Finally, we define $\rho$ denote the operator of multiplication by $c_{1}(T_{X})$.
@@
$\hat{\Gamma}(A)=\prod_{i=1}^{s}\Gamma(1+a_{i})$
where $a_{1},\ldots,a_{s}$ are the Chern roots of $A$, and $\Gamma(z)$ is the Gamma function expanded as a power series at $z=0$. Denote $\hat{\Gamma}(TX)$ by $\hat{\Gamma}_{X}$.
@@
$\Psi:K^{0}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
$A$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\right)$
@@
~$\widetilde{\mathcal{H}}_{X}=\mathcal{H}_{X}[\log z][z^{-1/2}].$
@@
(3.1.1) $\psi:\hat{H}^{*}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
$\phi$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\phi\right)$
@@
$\chi(A):=\sum_{i=0}^{\dim(X)}(-1)^{i}\operatorname{ch}^{\mathbb{T}}(H^{i}(X,A))$
where, as before, we use the equivariant Chern character with respect to $\mathbb{T}$. If $X$ is compact, then by [EG00, Tho92] we have the following equivariant Hirzebruch–Riemann–Roch theorem
(3.1.2) $\chi(A)=\int_{X}\operatorname{ch}^{\mathbb{T}}(A)\operatorname{Td}_{X},$
@@
$\int_{X}\alpha=\int_{Z}s^{*}(\alpha)/e_{\mathbb{T}}(V)$
where $s:Z\to V$ denotes the zero section. On the other hand, the projection formula in $K$-theory implies
$\chi(A)=\chi(s^{*}(A)/\wedge^{\bullet}(V^{\vee})).$
Applying the Hirzebruch–Riemann–Roch theorem (3.1.2) to $Z$ then implies (3.1.2) for $X$.
Following [CIJ18], consider a modified Euler characteristic $\chi_{z}$ given by
$\chi_{z}(C,D):=\left(\frac{2\pi\mathbbm{i}}{z}\right)^{\lambda\partial_{\lambda}}\chi(C,D)$
where $\lambda\partial_{\lambda}:=\sum_{i}\lambda_{i}\frac{\partial}{\partial\lambda_{i}}$ for some basis $\lambda_{1},\ldots,\lambda_{s}$ of $H^{2}(B\mathbb{T})$, .
Lemma 3.4 (Proposition 2.10 of [Iri09]).

### v2

@@ 2.3 Fourier–Mukai transform
$\overline{\mathbb{FM}}\colon\ \hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{-}\bigr{)}\to\hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{+}\bigr{)}$
denote the Fourier–Mukai transform in this case, where $\hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}$ denotes the completion of $K^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}$ along the augmentation ideal (see [19]).
By definition of the $\mathbb{T}$ action on $B$, $E$ and $F$, the (completion of the) equivariant cohomology of $B$ is given by
$\hat{H}^{*}_{\mathbb{T}}(B)=H^{*}(B)[\![z_{i},x_{i}]\!]_{i=1}^{n},$
where recall that $x_{i}=c_{1}^{\mathbb{T}}\bigl{(}L_{i}^{\vee}\bigr{)}$ and $z_{i}=c_{1}^{\mathbb{T}}\bigl{(}M_{i}^{\vee}\bigr{)}$. The equivariant line bundles $L_{i}\to B$ and $M_{i}\to B$ for $1\leq i\leq n$ determine a map
$v\colon\ \big{[}B/(\mathbb{C}^{*})^{2n}\big{]}\to\big{[}\{pt\}/(\mathbb{C}^{*})^{2n}\big{]}.$
@@
${\bigl{[}X_{\pm}/(\mathbb{C}^{*})^{2n}\bigr{]}}$${\bigl{[}\overline{X}_{\pm}/(\mathbb{C}^{*})^{2n}\bigr{]}}$${\bigl{[}B/(\mathbb{C}^{*})^{2n}\bigr{]}}$[${\bigl{[}\{pt\}/(\mathbb{C}^{*})^{2n}\bigr{]}.}$$\scriptstyle{\pi_{\pm}}$$\scriptstyle{\tilde{v}_{\pm}}$$\scriptstyle{\overline{\pi}_{\pm}}$$\scriptstyle{v}$
@@
with $H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\dots,\sigma_{r}]/\hat{I}_{\pm},$ we have the identification
${\pi_{\pm}}_{*}=\operatorname{id}_{B}\otimes_{\mathbb{C}}{\overline{\pi}_{\pm}}_{*},$
@@
$\hat{\operatorname{ch}}^{\mathbb{T}}\colon\ \hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}\to\hat{H}^{*}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}$
is an isomorphism (see [19]; here we use $\hat{\operatorname{ch}}^{\mathbb{T}}$ to denote the induced map defined on the completion $\hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{\pm}\bigr{)}$). Because it is induced by a derived equivalence, the map $\overline{\mathbb{FM}}$ is also an isomorphism. Proposition 2.10 implies that $\overline{\mathbb{U}}_{H}:=\hat{\operatorname{ch}}^{\mathbb{T}}\circ\overline{\mathbb{FM}}\circ\bigl{(}\hat{\operatorname{ch}}^{\mathbb{T}}\bigr{)}^{-1}$ is also an isomorphism, and may be defined before localizing with respect to $\mathbb{T}$. In other words, $\overline{\mathbb{U}}_{H}$ gives a well-defined map
$\overline{\mathbb{U}}_{H}\colon\ \mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\dots,\sigma_{r}]/\hat{I}_{-}\to\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\dots,\sigma_{r}]/\hat{I}_{+}.$
Again using the identification $\hat{H}^{*}_{\mathbb{T}}(X_{\pm})=H^{*}(B)\otimes_{\mathbb{C}}\mathbb{C}[\![z_{i},x_{i}]\!]_{i=1}^{n}[\sigma_{1},\dots,\sigma_{r}]/\hat{I}_{\pm}$ together with the explicit formula for $\mathbb{U}_{H}$ in (2.8), we also see that $\mathbb{U}_{H}=\operatorname{id}_{B}\otimes_{\mathbb{C}}\overline{\mathbb{U}}_{H}$ which implies the first statement.
Given $\overline{\alpha}\in\hat{H}^{*}_{\mathbb{T}}\bigl{(}\overline{X}_{-}\bigr{)}$, let $A=\bigl{(}\operatorname{ch}^{\mathbb{T}}\bigr{)}^{-1}(\overline{\alpha})\in\hat{K}^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{-}\bigr{)}$. Applying the equivariant Hirzebruch–Riemann–Roch theorem (see (3.2) below) twice, we see that for $A\in\hat{K}^{0}_{\mathbb{T}}(X_{-})$,
${\overline{\pi}_{-}}_{*}\bigl{(}\overline{\alpha}\operatorname{Td}_{\overline{X}_{-}}\bigr{)}$ $=\chi(A)=\chi\bigl{(}\overline{\mathbb{FM}}(A)\bigr{)}={\overline{\pi}_{+}}_{*}\bigl{(}\operatorname{ch}^{\mathbb{T}}\bigl{(}\overline{\mathbb{FM}}(A)\bigr{)}\operatorname{Td}_{\overline{X}_{+}}\bigr{)}$
$={\overline{\pi}_{+}}_{*}\bigl{(}\mathbb{U}_{H}(\overline{\alpha})\operatorname{Td}_{\overline{X}_{+}}\bigr{)}.$ (2.10)
The pullback $\tilde{v}_{\pm}^{*}\bigl{(}\operatorname{Td}_{\overline{X}_{\pm}}\bigr{)}$ is the relative Todd class $\operatorname{Td}_{X_{\pm}}$. This fact together with (2.10) allows us to upgrade (2.9) to the following commutative diagram:
@@
3 Gromov–Witten theory and $\bm{I}$-functions
Let $X$ be a smooth semi-projective variety. In what follows, $X$ will also be equipped with an action by a torus $\mathbb{T}$. Following [15], we assume that all $\mathbb{T}$-weights appearing in the representation $H^{0}(X,\mathcal{O}_{X})$ lie in a strictly convex cone, and that $H^{0}(X,\mathcal{O}_{X})^{\mathbb{T}}=\mathbb{C}.$ This implies in particular that $X^{\mathbb{T}}$ is compact.
@@ 3 Gromov–Witten theory and $\bm{I}$-functions
$\int_{X}\gamma:=\sum_{F}\int_{F}\frac{(\gamma)|_{F}}{e_{\mathbb{T}}\bigl{(}N_{F|X}\bigr{)}}\in S_{\mathbb{T}},$
@@
$\left\langle\alpha,\beta\right\rangle_{X}:=\sum_{F}\int_{F}\frac{(\alpha\cup\beta)|_{F}}{e_{\mathbb{T}}\bigl{(}N_{F|X}\bigr{)}}.$
@@
$R[\![Q]\!]=\biggl{\{}\sum_{d\in\operatorname{Eff}(X)}a_{d}Q^{d}\mid a_{d}\in R\biggr{\}}$
@@
$\mathcal{H}_{X}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\bigl{(}\!\bigl{(}z^{-1}\bigr{)}\!\bigr{)}[\![Q]\!]$
@@
$\Omega(f,g)=\operatorname{Res}_{z=0}\left\langle f(-z),g(z)\right\rangle_{X}{\rm d}z.$
@@
with $\mathcal{H}_{X}^{+}=H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}[z][\![Q]\!]$ and $\mathcal{H}_{X}^{-}=z^{-1}H^{*}_{\mathbb{T}}(X)\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\bigl{[}\!\bigl{[}z^{-1}\bigr{]}\!\bigr{]}[\![Q]\!]$.
For $\alpha_{1},\dots,\alpha_{m}\in H^{*}_{\mathbb{T}}(X)$, $k_{1},\dots,k_{m}\geq 0$, and $d\in\operatorname{Eff}(X)$, define the degree-$d$ genus-0 Gromov–Witten invariant of $X$:
$\big{\langle}\alpha_{1}\psi_{1}^{k_{1}},\dots,\alpha_{m}\psi_{m}^{k_{m}}\bigr{\rangle}_{0,m,d}^{X}=\int_{[\mathcal{M}_{0,m}(X,d)]^{vir}}\prod_{i=1}^{m}ev_{i}^{*}(\alpha_{i})\psi_{i}^{k_{i}}.$
If $X$ is not compact, the integral on the right-hand side is defined after localization.
@@
$\mathcal{F}_{X}({\bm{t}}(z))=\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle{\bm{t}}(\psi_{1}),\dots,{\bm{t}}(\psi_{m})\right\rangle_{0,m,d}^{X},$
@@
$\mathcal{L}_{X}=\bigl{\{}(p,q)\in T^{*}\mathcal{H}_{X}^{+}\mid p=d_{q}\mathcal{F}_{X}\bigr{\}}.$
@@
$-z1+{\bm{t}}(z)+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},{\bm{t}}(\psi_{2}),\dots,{\bm{t}}(\psi_{m+1})\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
@@
$J^{X}(t,z)=-z1+t+\sum_{d\in\operatorname{Eff}(X)}\sum_{m\geq 0}\frac{Q^{d}}{m!}\sum_{i\in I}\left\langle\frac{\phi_{i}}{-z-\psi_{1}},t,\dots,t\right\rangle_{0,m+1,d}^{X}\phi^{i}.$
@@
$J^{X}(\tau,z)+\sum_{i\in I}C^{i}(z)\frac{\partial}{\partial{t^{i}}}J^{X}(t,z)|_{t=\tau},$
@@
3.1 Integral lattice
As predicted by mirror symmetry, the genus zero Gromov–Witten theory ($A$-model) of a space $X$ should be equivalent to the Gauss–Manin connection ($B$-model) of a family of mirror manifolds $\check{X_{t}}$. The integral cohomology $H^{*}\bigl{(}\check{X_{t}},\mathbb{Z}\bigr{)}$ naturally endows the Gauss–Manin connection with an integral lattice of flat sections. Thus by mirror symmetry, one should expect an integral structure to appear in Gromov–Witten theory as well. Such a structure was defined by Iritani in [22] where it was verified to be compatible with mirror symmetry in the toric setting.
In this section, we describe Iritani’s integral structure on $\tilde{H}_{X}$. We define a lattice in $\tilde{H}_{X}$ as the image of $K_{\mathbb{T}}(X)$ under a certain modification of the Chern character. A key ingredient is a characteristic class known as the $\hat{\Gamma}$ class, which plays the role of a square root of the Todd class. This integral structure will allow us to compare the symplectic transformation $\mathbb{U}$ with the Fourier–Mukai transform $\mathbb{FM}$ from Section 2.3.
@@ 3.1 Integral lattice
$\mu(\phi)=\biggl{(}\frac{\deg(\phi)}{2}-\frac{\dim(X)}{2}\biggr{)}\phi,$
@@
Secondly, we define $\deg_{0}\colon H^{*}_{\mathbb{T}}(X)\to H^{*}_{\mathbb{T}}(X)$ be the degree operator, defined by
@@
Finally, we define $\rho$ to be the operator of multiplication by $c_{1}(T_{X})$.
@@
$\hat{\Gamma}(A)=\prod_{i=1}^{s}\Gamma(1+a_{i}),$
where $a_{1},\dots,a_{s}$ are the Chern roots of $A$, and $\Gamma(z)$ is the Gamma function expanded as a power series at $z=0$. Denote $\hat{\Gamma}(TX)$ by $\hat{\Gamma}_{X}$.
@@
$\Psi\colon\ K^{0}_{\mathbb{T}}(X)\to\widetilde{\mathcal{H}}_{X},\qquad A\mapsto z^{-\mu}z^{\rho}\bigl{(}\hat{\Gamma}_{X}\cup(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\bigr{)},$
@@
~$\widetilde{\mathcal{H}}_{X}=\mathcal{H}_{X}[\log z]\big{[}z^{-1/2}\big{]}.$
@@
$\psi\colon\ \hat{H}^{*}_{\mathbb{T}}(X)\to\widetilde{\mathcal{H}}_{X},\qquad\phi\mapsto z^{-\mu}z^{\rho}\bigl{(}\hat{\Gamma}_{X}\cup(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\phi\bigr{)}$ (3.1)
@@
$\chi(A):=\sum_{i=0}^{\dim(X)}(-1)^{i}\operatorname{ch}^{\mathbb{T}}\bigl{(}H^{i}(X,A)\bigr{)},$
where, as before, we use the equivariant Chern character with respect to $\mathbb{T}$. If $X$ is compact, then by [19, 32] we have the following equivariant Hirzebruch–Riemann–Roch theorem
$\chi(A)=\int_{X}\operatorname{ch}^{\mathbb{T}}(A)\operatorname{Td}_{X},$ (3.2)
@@
$\int_{X}\alpha=\int_{Z}s^{*}(\alpha)/e_{\mathbb{T}}(V),$
where $s\colon Z\to V$ denotes the zero section. On the other hand, the projection formula in $K$-theory implies
$\chi(A)=\chi\bigl{(}s^{*}(A)/\wedge^{\bullet}\bigl{(}V^{\vee}\bigr{)}\bigr{)}.$
Applying the Hirzebruch–Riemann–Roch theorem (3.2) to $Z$ then implies (3.2) for $X$.
Following [14], consider a modified Euler characteristic $\chi_{z}$ given by
$\chi_{z}(C,D):=\biggl{(}\frac{2\pi{\rm i}}{z}\biggr{)}^{\lambda\partial_{\lambda}}\chi(C,D),$
where $\lambda\partial_{\lambda}:=\sum_{i}\lambda_{i}\frac{\partial}{\partial\lambda_{i}}$ for some basis $\lambda_{1},\dots,\lambda_{s}$ of $H^{2}(B\mathbb{T})$.
Lemma 3.4 ([22, Proposition 2.10]).